# Euclidean lengths and the Culler-Shalen norms of slopes

### v1

@@ 2.3. Culler-Shalen norm
In this subsection, we define the Culler-Shalen norm of a slope and state a lemma used in the next section. The theory of such norms was originally developed by Culler-Shalen. See [22] for a survey. Also see [8] as a basic reference for this subsection.
@@
We denote by $R(\pi_{1}(M))$ the set of representations of $\pi_{1}(M)$ into $SL_{2}(\mathbb{C})$. It is naturally regarded as a complex affine algebraic set. We denote by $X(\pi_{1}(M))$ the set of all characters of representations in $R(\pi_{1}(M))$. This $X(\pi_{1}(M))$ is endowed with the structure of an affine algebraic set as in [6]. Thus it is called the character variety of $\pi_{1}(M)$. For $M$, there exists a particular $SL_{2}(\mathbb{C})$-representation induced from the hyperbolic structure. That is a discrete, faithful representation of $\pi_{1}(M)$ into $SL_{2}(\mathbb{C})$, called the holonomy representation. Please see [24] for a basic reference. Then we define a principal component $X_{0}$ of the character variety $X(\pi_{1}(M))$ to be a component that contains the character of the holonomy representation. This $X_{0}$ is known to be an irreducible complex affine algebraic curve due to Thurston. See [22] for detailed explanations. For $X_{0}$, we denote by $\tilde{X_{0}}$ the unique smooth projective curve that admits a birational correspondence $\phi:\tilde{X_{0}}\to X_{0}$. That is, the curve $\tilde{X_{0}}$ is constructed by desingularizing a projective completion of $X_{0}$. A point $x\in\tilde{X_{0}}$ is said to be an ideal point if it does not correspond to any point of $X_{0}$ under $\phi$. If $\gamma$ is an element of $\pi_{1}(M)$, then let $I_{\gamma}:X_{0}\to\mathbb{C}$ denote the rational function on $X_{0}$ defined by $I_{\gamma}(\chi)=\chi(\gamma)=\operatorname{trace}(\rho(\gamma))$.
@@
The next proposition assures the existence of a norm on $H_{1}(\partial M;\mathbb{R})$, which is now called the Culler-Shalen norm.
Proposition 2.4 (Culler-Shalen [8, Proposition 5.7]).
[…]
$\|\alpha\|=2\deg(I_{[c]}|_{X_{0}})$
holds, where $[c]$ denotes the element in $\pi_{1}(\partial M)$ given by a closed curve $c$ representing $\alpha$.
@@
Moreover, there are strict essential surfaces $S_{1},\cdots,S_{n}$ in $M$ corresponding to the ideal points $x_{1},\cdots,x_{m}$ such that
(1) $\|\alpha\|=\sum^{n}_{i=1}\Delta_{\partial M}(c,\partial S_{i})$
@@
holds for any homology class $\alpha\in H_{1}(\partial M;\mathbb{Z})\subset H_{1}(\partial M;\mathbb{R})$ represented by a closed curve $c$ in $\partial M$. Here $\Delta_{\partial M}(c,\partial S_{i})$ denotes the minimal geometric intersection number on $\partial M$ between the closed curve $c$ and $\partial S_{i}$, which we regard as the parallel family of the closed curves.
@@
Based on this proposition, we have the following lemma, which is well-known to specialists.
@@
Let $s_{1},\cdots,s_{n}$ be all the boundary slopes on $\partial M$. For any element $\boldsymbol{\alpha}\in H_{1}(\partial M;\mathbb{Z})\subset H_{1}(\partial M;\mathbb{R})$,
(2) $\|\alpha\|=\sum_{i=1}^{n}a_{i},\Delta(r,s_{i})$
holds, where each $a_{i}$ is a non-negative even integer, and $r$ is the slope corresponding to the element $\boldsymbol{\alpha}\in H_{1}(\partial M;\mathbb{Z})\subset H_{1}(\partial M;\mathbb{R})$. Moreover, if $s_{i}$ is associated to an ideal point, then $a_{i}$ is strictly positive.
@@
From equation (1), the equality (2) in the lemma is obtained by setting $a_{j}=0$ if $s_{j}$ does not correspond to $S_{i}$ in the proposition above.
Also, by [8, Proposition 3.10], each $S_{i}$ in the proposition is separating in $M$. This implies that $\Delta_{\partial M}(c,\partial S_{i})=a_{i},\Delta(r,s_{i})$ holds, where $a_{i}$ denotes the number of connected components of $\partial S_{i}$, which must be a non-negative even integer. ∎
As explained above, there is a natural two-to-one map from the set of primitive elements of $H_{1}(\partial M;\mathbb{Z})$ onto the set of slopes on $\partial M$. That is, a slope $r$ corresponds to ±$\pm\boldsymbol{r}\in H_{1}(\partial M;\mathbb{Z})$. Since $\|\boldsymbol{r}\|=\|-\boldsymbol{r}\|$ holds, we define $\|r\|$ by setting $\|r\|=\|\boldsymbol{r}\|=\|-\boldsymbol{r}\|$. In the rest of the paper, we will use this notation.
@@ 3. Lower bound on Culler-Shalen norm
Suppose that there exist two boundary slopes $s_{1},s_{2}$ on $\partial M$ such that they are associated with some ideal points of $X(\pi_{1}(M))$, and some essential surfaces $S_{1},S_{2}$ with boundary slopes $s_{1},s_{2}$ satisfy
@@
where $\chi_{i}$ denotes the Euler characteristic of $S_{i}$ and $b_{i}$ the number of boundary components of $S_{i}$ for $i=1,2$, respectively. Then
@@
Let $s_{1},\cdots,s_{n}$ be boundary slopes on $\partial M$. By Lemma 2.5, we have $\|r\|=\sum^{n}_{i=1}a_{i}\Delta(r,s_{i})$ for a slope $r$, with non-negative even integers $a_{1},\cdots,a_{n}$.
By assumption, there exist two boundary slopes, say $s_{1},s_{2}$, on $\partial M$ associated with some ideal points of $X(\pi_{1}(M))$. This implies that $a_{1}$ and $a_{2}$ are both strictly positive, in fact, they are both at least $2$, by Lemma 2.5.
@@
Let $M$ be the exterior of a hyperbolic two-bridge knot or the exterior of a $(-2,3,n)$-pretzel knot with $n$ odd and at least $7$ in $S^{3}$.
By virtue of Proposition 3.1, it suffices to show that there exist two boundary slopes $s_{1},s_{2}$ on $\partial M$ such that they are associated to some ideal points of $X(\pi_{1}(M))$, and some essential surfaces $S_{1},S_{2}$ with boundary slopes $s_{1},s_{2}$ satisfy
@@
First, consider the case where $M$ is the exterior of a hyperbolic two-bridge knot. Then it is proved in [20] that every boundary slope on $\partial M$ is associated to an ideal point of $X(\pi_{1}(M))$. On the other hand, we know that any two-bridge knot is alternating, i.e., it admits a diagram with alternatively arranged over-crossings and under-crossings running along it. Then, by [9], both checkerboard surfaces $S_{1}$ and $S_{2}$ for a reduced alternating diagram of an alternating knot are essential. Let $C$ be the number of crossings of the alternating diagram, which is actually the minimal crossing number of the knot ([15, 19, 23]). It is well-known that the distance $\Delta(s_{1},s_{2})$ of the boundary slopes $s_{1}$ and $s_{2}$ of $S_{1}$ and $S_{2}$ is equal to $2C$. Also, we see that $b_{i}=1$ for the number of boundary components $b_{i}$ of $S_{i}$ and
@@
Next, consider the case where $M$ is the exterior of a $(-2,3,n)$-pretzel knot with $n$ odd and at least $7$ in $S^{3}$. Then, by [12, 5.3.3. Corollaries], there are two boundary slopes $s_{1}=16$ and $s_{2}=2n+6$ of two essential surfaces $S_{1}$ and $S_{2}$ with the number of boundary components $b_{1}=b_{2}=1$ and the Euler characteristics $\chi_{1}=6-n$ and $\chi_{2}=1$, respectively. For these surfaces, we have the following as desired.
@@
The following statement is given as [8, Corollary 7.6]. Suppose that $M$ is a knot manifold (i.e., a compact, irreducible, orientable 3–manifold whose boundary is a torus) and that $M$ has at most two distinct isotopy classes of strict essential surfaces and $M$ is neither Seifert fibered nor an exceptional graph manifold. Let $S_{1}$ and $S_{2}$ be representatives of the two isotopy classes of connected strict essential surfaces. Let $s_{i}$ denote the boundary slope of $S_{i}$, $b_{i}$ denote the number of boundary components of $S_{i}$, and $\chi_{i}$ denote the Euler characteristic of $S_{i}$ for $i=1,2$. Then for $i=1,2$ we have
@@
In Proposition 3.1, the first condition that the two boundary slopes $s_{1},s_{2}$ on $\partial M$ are associated to some ideal points of $X(\pi_{1}(M))$ is necessary for our proof, but it is not easy to verify for a given knot. On the other hand, the second condition on the slopes $s_{1},s_{2}$ can be satisfied for large classes of knots, as described below.
@@
Suppose next that $M$ is the exterior of a hyperbolic Montesinos knot $K$. Then, the following holds by [13, Corollary 1.4]. Among its non-meridional boundary slopes, let $s_{1}$ and $s_{2}$ be the maximum and minimum, respectively. Then, there exist two essential surfaces $S_{1}$ and $S_{2}$ with boundary slopes $s_{1}$ and $s_{2}$ such that
@@
For example, consider the exterior $M$ of a $(-2,3,n)$-pretzel knot in $S^{3}$ with $n$ odd, at least $7$, and $n\not\equiv 0\pmod{3}$. Then, by [17], the norm $\|m\|$ of the meridional slope $m$ is equal to $93n-$ 9. However, since Dehn filling along $m$ yields $S^{3}$, which is exceptional, the Euclidean length of the slope $m$ is at most 6 by [2, 16]. ∎
@@ 4.1. Lengths of numerical slopes
Let $r_{1}=p_{1}/q_{1}$ and $r_{2}=p_{2}/q_{2}$ be non-meridional numerical slopes on $\partial M$. We choose a horotorus $T$ and identify $\partial M$ with $T$. Let $\widetilde{T}$ be a component of the preimage of $T$ in the universal cover of Int$(M)$. This $\widetilde{T}$ is naturally identified with the Euclidean plane $\mathbb{E}^{2}$. The preimage of a point on $T$ gives a lattice on $\widetilde{T}$ with base point $O$ as utilized in [10]. Then, if a primitive lattice point $X$ corresponds to a slope $r$ on $T$, then the distance $d(O,X)$ between $O$ and $X$ equals the length $\mathrm{length}_{T}(r)$ of the slope $r$.

### v2

@@ 2.3. Culler-Shalen norm
In this subsection, we define the Culler–Shalen norm of a slope and state a lemma used in the next section. The theory of such norms was originally developed by Culler and Shalen. See [10] as a basic reference for this subsection. See also [24] for a survey.
@@
We denote by $R(\pi_{1}(M))$ the set of representations of $\pi_{1}(M)$ into $SL_{2}(\mathbb{C})$. It is naturally regarded as a complex affine algebraic set. We denote by $X(M)$ the set of all characters of representations in $R(\pi_{1}(M))$. The set $X(M)$ is endowed with the structure of an affine algebraic set as in [8]. Thus it is called the character variety of $\pi_{1}(M)$, or simply of $M$. For a hyperbolic 3-manifold $M$, there exists a particular $SL_{2}(\mathbb{C})$-representation of $\pi_{1}(M)$ induced from the hyperbolic structure. This is a discrete, faithful representation of $\pi_{1}(M)$ into $SL_{2}(\mathbb{C})$, the so-called holonomy representation. See [26] for a basic reference. Then we define a principal component $X_{0}$ of the character variety $X(M)$ to be a component that contains the character of the holonomy representation. This $X_{0}$ is known to be an irreducible one-dimensional component of $X(M)$, i.e., a complex affine algebraic curve, due to Thurston. See [24] for detailed explanations. For $X_{0}$, we denote by $\tilde{X_{0}}$ the unique smooth projective curve that admits a birational correspondence $\phi:\tilde{X_{0}}\to X_{0}$. That is, the curve $\tilde{X_{0}}$ is constructed by desingularization to obtain a projective completion of $X_{0}$. A point $x\in\tilde{X_{0}}$ is said to be an ideal point if it does not correspond to any point of $X_{0}$ under $\phi$. If $\gamma$ is an element of $\pi_{1}(M)$, then let $I_{\gamma}:X_{0}\to\mathbb{C}$ denote the rational function on $X_{0}$ defined by $I_{\gamma}(\chi)=\chi(\gamma)=\operatorname{trace}(\rho(\gamma))$.
@@
The next proposition assures the existence of a norm on $H_{1}(\partial M;\mathbb{R})$, which is called the Culler-Shalen norm associated to $X_{0}$.
Proposition 2.4 (Culler–Shalen [10, Proposition 5.7]).
Let $X_{0}$ be a principal component of $X(M)$ and let $x_{1},\cdots,x_{n}$ denote the ideal points of $\tilde{X_{0}}$. Then there exists a unique norm $\|\cdot\|_{0}$ on the vector space $H_{1}(\partial M;\mathbb{R})$ such that for any element $\alpha\in H_{1}(\partial M;\mathbb{Z})\subset H_{1}(\partial M;\mathbb{R})$, it satisfies
$\|\alpha\|_{0}=\deg\!\left({I_{[c]}|_{X_{0}}}^{2}-4\right)=2\,\deg\!\left(I_{[c]}|_{X_{0}}\right).$
@@
Here $[c]$ denotes the element in $\pi_{1}(\partial M)$ represented by a closed curve $c$ corresponding to $\alpha$. Moreover, there are strict essential surfaces $S_{1},\cdots,S_{n}$ in $M$ corresponding to the ideal points $x_{1},\cdots,x_{n}$ of $\tilde{X_{0}}$ such that
(1) $\|\alpha\|_{0}=\sum_{j=1}^{n}\Delta_{\partial M}(c,\partial S_{j})$
@@
holds for any homology class $\alpha\in H_{1}(\partial M;\mathbb{Z})\subset H_{1}(\partial M;\mathbb{R})$ represented by a closed curve $c$ in $\partial M$. Here $\Delta_{\partial M}(c,\partial S_{j})$ denotes the minimal geometric intersection number on $\partial M$ between the closed curve $c$ and $\partial S_{j}$, which we regard as a parallel family of closed curves.
We remark that each $S_{j}$ in the proposition above is the surface captured by the action on the tree $T_{x_{j}}$ associated to the ideal point $x_{j}$, which is called a $T_{x_{j}}$-surface in [10]. In fact, it is a possibly disconnected, separating surface properly embedded in $M$. See [10] for further details.
The same machinery as above almost works for any one-dimensional component of the character variety $X(M)$ that contains the character of an irreducible representation, and one can, in general, obtain a seminorm associated with the component. Then, the (total) Culler-Shalen norm $\|\cdot\|$ on $H_{1}(\partial M;\mathbb{R})$ is defined as the total sum of the values of the seminorms on the irreducible one-dimensional components of $X(M)$.
If the seminorm obtained for a one-dimensional component (curve) of $X(M)$ is a genuine norm, then the component is called a norm curve. By Proposition 2.4, the principal component $X_{0}$ is a norm curve.
@@
Based on these settings, we have the following lemma, which is well known to specialists.
@@
Let $s_{1},\dots,s_{m}$ be all the boundary slopes on $\partial M$. For any element $\boldsymbol{\alpha}\in H_{1}(\partial M;\mathbb{Z})\subset H_{1}(\partial M;\mathbb{R})$,
(2) $\|\boldsymbol{\alpha}\|=\sum_{i=1}^{m}a_{i}\,\Delta(r,s_{i})$
holds, where each $a_{i}$ is a non-negative integer, and $r$ is the slope corresponding to $\boldsymbol{\alpha}\in H_{1}(\partial M;\mathbb{Z})\subset H_{1}(\partial M;\mathbb{R})$. Moreover, if $s_{i}$ is associated to an ideal point of some norm curve of $X(M)$, then $a_{i}$ is a positive even integer.
@@
Consider first the Culler-Shalen norm $\|\cdot\|_{0}$ associated with a principal component $X_{0}$. From Equation (1), the equality for $\|\cdot\|_{0}$ in the lemma is obtained by setting $a_{i}=0$ whenever $s_{i}$ does not correspond to a surface $S_{j}$ as in the proposition above. Also, by [10, Proposition 3.10], each $S_{j}$ in the proposition is a separating surface in $M$. This implies that, by setting $a_{i}$ as the sum of the numbers of connected components of the $\partial S_{j}$’s, the sum of $\Delta_{\partial M}(c,\partial S_{j})$’s for the surfaces $S_{j}$ with boundary slope $s_{i}$ is equal to $a_{i}\,\Delta(r,s_{i})$. Since $\Delta_{\partial M}(c,\partial S_{j})$ denotes the minimal geometric intersection number between $c$ and $\partial S_{j}$, and $S_{j}$ is a (possibly disconnected) separating surface, each $a_{i}$ must be a positive even integer. For the total Culler-Shalen norm $\|\cdot\|$, by definition, the same arguments imply the statements in the lemma. ∎
As explained above, there is a natural two-to-one map from the set of primitive elements of $H_{1}(\partial M;\mathbb{Z})$ onto the set of slopes on $\partial M$. That is, a slope $r$ corresponds to $\pm\boldsymbol{r}\in H_{1}(\partial M;\mathbb{Z})$. Since $\|\boldsymbol{r}\|=\|-\boldsymbol{r}\|$ holds, we define $\|r\|$ for a slope $r$ by setting $\|r\|=\|\boldsymbol{r}\|=\|-\boldsymbol{r}\|$. In the rest of the paper, we will use this notation.
@@ 3. Lower bound on Culler-Shalen norm
Suppose that there exist two boundary slopes $s_{1},s_{2}$ of essential surfaces $S_{1},S_{2}$ in $M$ such that they are associated with distinct ideal points of some norm curves in $X(M)$. Let $\chi_{i}$ denote the Euler characteristic of $S_{i}$ and $b_{i}$ the number of boundary components of $S_{i}$ for $i=1,2$, respectively. If the inequality
@@
is satisfied, then
@@
Let $s_{1},\cdots,s_{n}$ be boundary slopes on $\partial M$. By Lemma 2.5, we have $\|r\|=\sum^{n}_{i=1}a_{i}\Delta(r,s_{i})$ for a slope $r$, with non-negative integers $a_{1},\cdots,a_{n}$ with $a_{i}\geq 2$.
By assumption, there exist two boundary slopes, say $s_{1},s_{2}$, on $\partial M$ associated with some ideal points of norm curves in $X(M)$. This implies that $a_{1}$ and $a_{2}$ are both strictly positive, in fact, they are both at least $2$, by Lemma 2.5.
@@
Let $M$ be the exterior of a hyperbolic two-bridge knot or the exterior of a (2,3,n)(-2,3,n)-pretzel knot with $n$ odd and at least $7$ in $S^{3}$.
By virtue of Proposition 3.1, it suffices to show that there exist two boundary slopes $s_{1},s_{2}$ on $\partial M$ such that they are associated to some ideal points of $X(M)$, and some essential surfaces $S_{1},S_{2}$ with boundary slopes $s_{1},s_{2}$ satisfy
@@
First, consider the case where $M$ is the exterior of a hyperbolic two-bridge knot. Then it is proved in [22] that every boundary slope on $\partial M$ is associated to an ideal point of norm curves in $X(M)$. See [4] for more explanations. In particular, as is stated in [16, 4.2], if the knot is twist knots, the character variety has only one norm curve $X_{0}$, and so $\|\cdot\|=\|\cdot\|_{0}$ holds. Also see [5] for this case. On the other hand, we know that any two-bridge knot is alternating, i.e., it admits a diagram with alternatively arranged over-crossings and under-crossings running along it. Then, by [11], both checkerboard surfaces $S_{1}$ and $S_{2}$ for a reduced alternating diagram of an alternating knot are essential. Let $C$ be the number of crossings of the alternating diagram, which is actually the minimal crossing number of the knot ([17, 21, 25]). It is well-known that the distance $\Delta(s_{1},s_{2})$ of the boundary slopes $s_{1}$ and $s_{2}$ of $S_{1}$ and $S_{2}$ is equal to $2C$. Also, we see that $b_{i}=1$ for the number of boundary components $b_{i}$ of $S_{i}$ and
@@
Next, consider the case where $M$ is the exterior of a (2,3,n)(-2,3,n)-pretzel knot with $n$ odd and at least $7$ in $S^{3}$. These are well-known to be hyperbolic, and the following are proved in [19]: The character variety $X(M)$ has only one norm curve $X_{0}$, and so $\|\cdot\|=\|\cdot\|_{0}$ holds. Moreover, every boundary slope on $\partial M$ is associated to an ideal point of the norm curve $X_{0}$.
On the other hand, by [14, 5.3.3. Corollaries], there are two boundary slopes $s_{1}=16$ and $s_{2}=2n+6$ of two essential surfaces $S_{1}$ and $S_{2}$ with the number of boundary components $b_{1}=b_{2}=1$ and the Euler characteristics $\chi_{1}=6-n$ and $\chi_{2}=1$, respectively. For these surfaces, we have the following as desired.
@@
The following statement is given as [10, Corollary 7.6]. Suppose that $M$ is a knot manifold (i.e., a compact, irreducible, orientable 3–manifold whose boundary is a torus) and that $M$ has at most two distinct isotopy classes of strict essential surfaces and $M$ is neither Seifert fibered nor an exceptional graph manifold. Let $S_{1}$ and $S_{2}$ be representatives of the two isotopy classes of connected strict essential surfaces. Let $s_{i}$ denote the boundary slope of $S_{i}$, $b_{i}$ denote the number of boundary components of $S_{i}$, and $\chi_{i}$ denote the Euler characteristic of $S_{i}$ for $i=1,2$. Then for $i=1,2$ we have
@@
In Proposition 3.1, the first condition that the two boundary slopes $s_{1},s_{2}$ on $\partial M$ are associated to some ideal points of norm curves in $X(M)$ is necessary for our proof, but it is not easy to verify for a given knot. On the other hand, the second condition on the slopes $s_{1},s_{2}$ can be satisfied for large classes of knots, as described below.
@@
Suppose next that $M$ is the exterior of a hyperbolic Montesinos knot $K$. Then, the following holds by [15, Corollary 1.4]. Among its non-meridional boundary slopes, let $s_{1}$ and $s_{2}$ be the maximum and minimum, respectively. Then, there exist two essential surfaces $S_{1}$ and $S_{2}$ with boundary slopes $s_{1}$ and $s_{2}$ such that
@@
For example, consider the exterior $M$ of a (2,3,n)(-2,3,n)-pretzel knot in $S^{3}$ with $n$ odd, at least $7$, and $n\not\equiv 0\pmod{3}$. Then, by [19], the norm $\|m\|$ of the meridional slope $m$ is equal to 3n93n-9. However, since Dehn filling along $m$ yields $S^{3}$, which is exceptional, the Euclidean length of the slope $m$ is at most 6 by [2, 18]. ∎
@@ 4.1. Lengths of numerical slopes
Let $r_{1}=p_{1}/q_{1}$ and $r_{2}=p_{2}/q_{2}$ be non-meridional numerical slopes on $\partial M$. We choose a horotorus $T$ and identify $\partial M$ with $T$. Let $\widetilde{T}$ be a component of the preimage of $T$ in the universal cover of Int(M)(M). This $\widetilde{T}$ is naturally identified with the Euclidean plane $\mathbb{E}^{2}$. The preimage of a point on $T$ gives a lattice on $\widetilde{T}$ with base point $O$ as utilized in [12]. Then, if a primitive lattice point $X$ corresponds to a slope $r$ on $T$, then the distance $d(O,X)$ between $O$ and $X$ equals the length $\mathrm{length}_{T}(r)$ of the slope $r$.